OpenStudy (anonymous):

A charged particle of mass 'm' and having a charge 'Q' is placed in an electric field E which varies with time as E = E' sin (wt) ----(w=omega) what is the amplitude of SHM executed by the particle? a) QE'^2/mw^2 b)Q E'/mw^2 c)sqrt(QE'/mw^2) d)QE'/mw

OpenStudy (anonymous):

What is the general equation for a SHM @goutham1995 ??

OpenStudy (anonymous):

X = a sin(wt + theta)

OpenStudy (anonymous):

@goutham1995 , It should be logically option B .The simple approach should be Let \[\large x= \frac{QE'}{m {\omega}^2}\] \[\large m {\omega}^2x=QE'\] It is of the form \[\large m(a)= F\] a = -w^2x F = qE'

OpenStudy (anonymous):

Rest of options won't give this. :) So only option B is correct

OpenStudy (anonymous):

@goutham1995

OpenStudy (anonymous):

but how do you do it without taking those options?